OpenStudy (anonymous):

0.08 divided by 3575

OpenStudy (anonymous):

0.08 / 3575

OpenStudy (anonymous):

2.442748092E-5

OpenStudy (anonymous):

you cannt work with it as a decimal so what you do is convert it to a fractiona nd then divide

OpenStudy (anonymous):

ignore him he just trolled you lol

OpenStudy (anonymous):

yep so show me wat u think to do

OpenStudy (anonymous):

well i think you have to make 0.08 an whole number so divide it by 100 or move the decimal point two places to the right @HawkCrimson

OpenStudy (anonymous):

yep so show me it in an equation form

OpenStudy (anonymous):

can't

OpenStudy (anonymous):

.08 / 3575 will become (8/100) / 3575 can you continue it?

OpenStudy (anonymous):

help please @robtobey

OpenStudy (anonymous):

@hero

OpenStudy (anonymous):

@cwrw238 help please

OpenStudy (anonymous):

\[\frac{.08}{3575}=\frac{8}{357500}=0.0000223776=2.23776*10^{-5} \]